Note: Draw the P-V, T-S diagrams for all problems

13. The work required to compress a gas reversibly according to PV 1.3 =C is 68KJ. if there is no flow

determine the change in internal energy, enthalpy and heat if the gas is CO 2 .


14. Air is compressed polytropically so that the quantity PV 1.4 is constant. If 0.02 m 3 of air at atmospheric pressure of 101.3 KPA and 4C are compressed to a gage pressure of 405KPAg, determine the final temperature of air in C.


1
Expert's answer
2021-04-01T01:08:42-0400

13. pV1.3 = C

Q = 0

∆U = Q - W = -W

∆H = ∆U = -68 000 J

14. pV1.4 = const

V1 = 0.02 m3; p1 = 101.3 kPa; T1 = 4°C = 4 + 273 = 277 K;

p2 = 101.3 + 405 = 506.3 kPa

p1V1n = p2V2n

p1/p2 = (V2/V1)n

101.3/506.3 = (V2/0.02)1.4

V2 = 0.006 m3

Substance is ideal gas pV = RT

T = pV/R = 506.3 x 103 Pa x 0.006 m3 / 8.31 m3 Pa K-1 = 365 K